If a coin was tossed 1000 times what would be the probability that coin will land on heads?
1/2 or 0.5
Assuming they're numbered the way you're implying, you have a 2 out of 3 chance (2/3, 6/9) of getting an even tile. That's 66.66666666etc
1/6,3/6
-----------------------------------------------------------------------------------------------------
The probability of rolling a 2 is: P(2) = 1/6
The probability of rolling an even number is: P(even) = 1/2
The result on the second roll is independent of the result in the first roll.
The probability of rolling a 2 and then rolling an even number is:
P(2,even no.) = (1/6) ∙ (1/2) = 1/12 = 0.08333... ≈ 8.33%
If the order of choice does not matter, the expression is 25 X 24 X 23 X 22 X 21 X 20 X 19 X 18 X 17, which can also be written as 25!/16!.
What is the value of z if the area between -z and z is 0.754?
Since the normal distribution is symmetric, the area between -z and 0 must be the same as the area between 0 and z. Using this fact, you can simplify this problem to finding a z such that the area between 0 and z is .754/2=.377. If you look this value up in a z-table or use the invNorm on a calculator, you will find that the required value of z will be 1.16. Therefore, the area between -1.16 and 1.16 must be approximately .754.
What is the significance of the mean of a probability distribution?
The significance of the mean of a probability distribution is that it is the most probably thing to happen. The mean is the average of a set of values. If it is the average of a probability distribution, it is the most probable part.
8:6
What is the standard deviation of the scores 100 92 94 95 81 82 87 89 71 71 73 61 62 68 51 55?
The standard deviation of the scores 100 92 94 95 81 82 87 89 71 71 73 61 62 68 51 55 is: 15.1921
If you flip 5 coins how many possible outcomes are there?
If each coin is a different color, then there are 32 possible outcomes.
If you can't tell the difference between the coins, and you're just counting
the number of heads and tails, then there are 6 possible outcomes:
5 heads
4 heads
3 heads
2 heads
1 heads
all tails
2
Rolling a single die 34 times keeping track of the numbers rolled is this a bionomial distribution?
No. If you keep track of all the numbers that are rolled then there are six possible outcomes, a binomial has only two. It is a multinomial distribution.
If a fair coin is tossed eight times what is the probability of getting exactly 6 heads?
You need a formula for this.
If the probability (in one toss) of getting head is "p", then the probability of getting exactly k heads out of n tosses is:
(n,k) p^k (1-p)^(n-k)
where (n,k) denotes the number of combinations of k elements among n.
You should also know that (n,k) = n! / (( n-k)! k! )
so here, with n=8, k=6, and p=.5 you have
(n,k) = 8*7 / 2 = 28
and your probability is :
28 * 1/2^6 * 1/2^2 = 28 / 256 = 7 / 64
3.92
If the probability of a defect is 0.015 what is the probability of not having a defect?
1-.015 = .985
n event A is said to be less probable than an event B if the probability of A occurring is smaller than that of B occurring.
Why is the sample mean an unbiased estimator of the population mean?
The sample mean is an unbiased estimator of the population mean because the average of all the possible sample means of size n is equal to the population mean.
Suppose probability of drawing a red marble is p.
Then p = 2*(1 - p)
that is p = 2 - 2p
or p = 2/3
So 2/3 of the 24 marbles are red
24*(2/3) = 16 red marbles.
What is the probability of obtaining 3 heads in three flips of a fair coin?
1/2 * 1/2 * 1/2 = 1/8 = 12.5%